OpenStudy (anonymous):

ennifer has been saving for college for 57 months. The first month, she saved $11. She was able to save more money each month than the month before. She ended up saving $19,779.00. How much more did she save each month? $11.00 $11.50 $12.00 $12.50

OpenStudy (anonymous):

PLEASE HELP ILL GIVE MEDALS

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

so first month we saved 11 dollars how much did she saved the next month?

OpenStudy (anonymous):

idk im very confused

OpenStudy (freckles):

is that the whole question? it doesn't say how much was saved each month

OpenStudy (freckles):

or state a relationship between consecutive months

OpenStudy (anonymous):

ya its the whole question

OpenStudy (freckles):

There doesn't seem to be enough information. There is no relationship given between the months...

OpenStudy (freckles):

we could try to assume we have an arithmetic series and that the difference between each consecutive month is constant

OpenStudy (freckles):

but the question doesn't say to assume this

OpenStudy (freckles):

11,11+d,11+2d,...,11+nd \[\sum_{i=0}^{n}(11+di)=19779\] there are 57 months so n actually equals 57 \[\sum_{i=0}^{57}(11+di)=19779\]

OpenStudy (freckles):

but there is no real relationship it just saved Jennifer saved more and more which doesn't imply there is a common difference